when Tracy McCoy wakes up Saturday morning, she remembers

that she promised the PTA she would make some cakes and/or homemade bread for its bake sale

that afternoon. However, she does not have time to go to the store to get ingredients, and she has

only a short time to bake things in her oven. Because cakes and breads require different baking

temperatures, she cannot bake them simultaneously, and she has only 3 hours available to bake.

A cake requires 3 cups of flour, and a loaf of bread requires 8 cups; Tracy has 20 cups of flour.

A cake requires 45 minutes to bake, and a loaf of bread requires 30 minutes. The PTA will sell a

cake for $10 and a loaf of bread for $6. Tracy wants to decide how many cakes and loaves of

bread she should make.

a. Formulate a linear programming model for this problem.

b. Solve this model by using graphical analysis.


1
Expert's answer
2021-03-25T17:52:04-0400

Solution:

Let the number of cakes be "x" and number of loaves of bread be "y" .

Objective function: To maximise cost, "Z=10x+6y"

Subject to the constraints:

"3x+8y\le20
\\ 45x+30y\le3(60)\Rightarrow3x+2y\le12
\\x,y\ge0"

Consider them equations and plotting on the graph, we get following graph.

Also, put (0,0) in these inequations.

"0\le20\Rightarrow True
\\0\le12\Rightarrow True"

Thus, their shadow or shaded area is towards (0,0).



Now, OABC is the feasible region.

Solving these equations to get point B.

"O(0,0),A(4,0),B(\frac {28}9,\frac 43),C(0,2.5)"

At "O(0,0): Z=0+0=0"

At "A(4,0):Z=10(4)+0=40"

At "B(\frac {28}9,\frac 43):Z=10(\frac {28}9)+6(\frac 43)\approx31.1+12=43.1"

At "C(0,2.5):Z=0+6(2.5)=15"

Clearly, maximum is 43.1 at B.

Hence, she should make "\frac {28}9" cakes and "\frac 43" loaves of bread.